# Coherent estimation of risk measures

### v1

@@ Abstract
risk estimator , coherent risk estimator , estimation of risk measures , risk bias , estimation of risk , expected shortfall , value at risk , coherent risk measure , asymptotic properties of risk estimators , Basel framework , FRTB , IMA , market risk , regulatory capital model
@@ 1 Introduction
In this work, we focus on the second step and adopt a novel perspective, fundamentally different from the existing literature, focused on the estimation function of $\rho$, say $\hat{\rho}$, which is later used to estimate $\hat{\rho}(X)$. We argue that, similar to the risk measures themselves, any estimation must also satisfy a set of desirable financial normative properties, postulated a priori, and in addition be a ”good approximation” [^1] There is a subtle distinction between estimation and approximation. Estimation refers to inferring an unknown true value from limited data, with an emphasis on its statistical properties. Approximation, in contrast, involves simplifying a known value or formula to make it computationally tractable, while analyzing the resulting numerical error. Our approach in this work combines elements of both, but for consistency we refer to them jointly as estimation.. This approach is motivated by the recognition that risk quantification procedures serve primarily to determine capital reserves for mitigating exposures, rather than to provide mere approximations of intrinsic risk values. To this end, we introduce the notion of the coherent risk estimator (CRE) that maps actual P&L samples to real numbers, that is monotone, cash additive, positive homogeneous, and subadditive when applied to a sample X.
@@
Our key theoretical contribution is the development of the robust representations of all CREs, see Theorem 4.1. Similar results are obtained for law-invariant [^2] Similar to risk measures, a law-invariant CRE does not depend on the ordering of the input sample. See Definition 3.2. CREs, a sublcass of CREs. In fact, we prove that a CRE is law-invariant if and only if it can be represented as a supremum over a set of $L$-estimators; see Theorem 4.2. Moreover, assuming additionally that a CRE is also comonotonic, we show that it can be represented as an $L$-estimator; see Theorem 4.10. The importance of such results is evident. This approach provides a canonical framework for constructing risk estimators that are designed to possess the desired risk management properties. Furthermore, it enables not only the systematic selection of favorable estimators from the wide variety of existing alternatives, but also ensures that the chosen methods remain practically applicable, are integral, robust and distribution-free, and aligned with supervisory expectations, see e.g. ECB [2025]. Additionally, the strong connection between CREs and $L$-estimators facilitates the leverage of the extensive statistical literature about $L$-estimators and their various asymptotic properties, mostly applied to comonotonic and law-invariant CREs. Our approach also provides a fresh look on the robust representation theory for risk measures, in which the duality is build directly into the estimation formula rather than being imposed only on a theoretical risk metric level. Throughout the paper, we present various examples of CREs, with particular emphasis on estimators of $\operatorname{\mathrm{ES}}$ and their properties, including a dedicated section comparing different $\operatorname{\mathrm{ES}}$ estimators (see Section 6). We also discuss methods for constructing CREs for a given coherent risk measure using plug-in procedures, which recover estimators of several classical risk measures, including $\operatorname{\mathrm{ES}}$.
@@
Third, in the regulatory Internal Model Approach (IMA) for Pillar I bank models, the 10-day $\operatorname{\mathrm{VaR}}$ and 10-day $\operatorname{\mathrm{ES}}$ at the confidence levels $\alpha=1\%$ and $\alpha=2.5\%$, respectively, are the key reference market risk capital metrics, see BCBS [2006, 2019] for details. [^3] In most practical applications the confidence level $\alpha$ is set to $1\%$, $2.5\%$, or $5\%$. Two main conventions are used when quoting confidence levels for $\operatorname{\mathrm{VaR}}$ and $\operatorname{\mathrm{ES}}$: the left-tail convention, adopted in this work and common in the risk measurement literature, and the right-tail convention, which reports thresholds of the form $1-\alpha$ (e.g., $99\%$, $97.5\%$, $95\%$) and is more widespread in risk management and regulatory practice, where the right tail represents losses. Although the risk measures themselves are fixed, regulatory and supervisory bodies rarely prescribe explicit estimation formulas. Two notable exceptions are the following locally implemented formulas used for Pillar I capital calculations: the minimal Stressed VaR formula under the Basel II framework [PRA, 2020, Article 10.2] and the EU Stress Scenario Risk Measure under the Basel III framework [EU, 2024b, Article 11]. More generally, regulatory and supervisory texts specify desired properties of estimators – such as conceptual soundness, proven backtesting track record, or distribution-free character – rather than their explicit form, see e.g. [ECB, 2025, Section 5.3], [PRA, 2020, Section 10], [CBUAE, 2023, Market Risk Standards & Risk Management Standards], or [HKMA, 2024, Section 4.5]. In practice, market risk estimators are often non-parametric and based on order statistics from historical simulation P&Ls, i.e., sorted P&L sample values. For $\operatorname{\mathrm{VaR}}$ and $\operatorname{\mathrm{ES}}$, the estimators are typically linear combinations of order statistics, with coefficients independent of the distribution of $X$; see EBA [2025] for typical VaR look-back period and weighting choices. These observations further motivates the relevance of the (comonotonic) CRE representations developed in this paper. We refer to Section 6 for a more detailed discussion of various nonparametric $\operatorname{\mathrm{ES}}$ estimations.
@@ 3 Coherent risk estimators
    Monotonicity, for any $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{n}$ such that [^4] For vector order $\mathbf{x}\leq\mathbf{y}$ we use the component wise comparison $x_{i}\leq y_{i}$, for $i=1,2,\ldots,n$. $\mathbf{x}\geq\mathbf{x}^{\prime}$, we have $\hat{\rho}_{n}(\mathbf{x})\leq\hat{\rho}_{n}(\mathbf{x}^{\prime})$;
@@ 4.1 Comonotonic CREs and their representation as L-estimators
where $\xrightarrow{a.s.}$ stands for $\mathbb{P}$-almost sure convergence [^5] In this work, for simplicity, we focus on $\mathbb{P}$-a.s. convergence of the estimators, which corresponds to strong consistency in classical statistics. Nevertheless, most of the results can be extended to weaker forms of convergence, such as convergence in probability (i.e., weak consistency)..
@@ 6 Numerical study: comparison of ES estimators based on L-statistics
Taking into account the nature of the considered metrics, we categorise the first two metrics (AE and SE) as Fit metrics, the next two (SB and RB) as Bias metrics, and the last one (CT) as Confidence metric. It should be noted that while our assessment is focused on estimators’ statistical properties, ES estimators are effectively point-forecast metrics focused on tail risk quantification. Thus, due to the risk-oriented nature of the ES estimators, the output of the AE and SE fit metrics should be handled with care. Those metrics may not be optimal for comparative accuracy assessment (and selection algorithms), since they are not elicitable for ES; see Gneiting [2011]. In particular, ES itself is not an elicitable risk measure, which makes comparative estimator analysis based on backtesting or scoring challenging. It is known that the joint metric $(\operatorname{\mathrm{VaR}},\operatorname{\mathrm{ES}})$ is elicitable, see Fissler and Ziegel [2016], but the associated scoring functions are difficult to compute and interpret. Moreover, the resulting comparisons are often statistically insignificant, may depend on the specific choice of scoring function, and require joint consideration of both ES and VaR estimation methodologies. Due to these challenges, we have decided to present, in addition to AR and SE, another set of metrics related to bias, as well as confidence measurement. We recall that the purpose of this paper is to provide a comparison of ES estimators based on $L$-statistics in the context of their CRE representation and the chosen scheme. A comprehensive assessment and comparative performance analysis is beyond the scope of this paper. For in-depth quantitative studies based on both simulated and market data, we refer to Righi and Ceretta [2015] and Chen [2008].
@@
In the next numerical example, we calculate the benchmark metrics for selected families of distributions using Monte Carlo simulations. For comparability, we consider the same family of distributions as in [EBA, 2023, Annex I]. This constitutes eight different distributions: standard normal distribution, Student’s $t$-distribution with $\nu=5$ degrees of freedom, and six different normal inverse gamma (NIG) distributions with parameters $(\alpha,\beta)\in\{(0.4,0.14),(0.4,-0.14),(0.55,0.3025),(0.55,-0.3025),(0.4,0.22),(0.4,-0.22)\}$; the location parameter $\mu=0$ and the scale parameter $\delta=1$ are standardized. Skewness and kurtosis of the NIG distributions are known analytically and as NIG distributions are closed under convolution also for their rolling sums, [Scott et al., 2011, Section 6.1]. For the considered values of $(\alpha,\beta)$ parameters, the skewness and excess kurtosis values are $(0.9460,3.6282)$, (0.9460,3.6282)(-0.9460,3.6282), $(1.3601,4.5051)$, (1.3601,4.5051)(-1.3601,4.5051), $(1.8702,8.5174)$, and (1.8702,8.5174)(-1.8702,8.5174). The choice of $\alpha$ and $\beta$ aims to consider various signs and magnitudes of the skewness and the kurtosis comparable to trading book P&Ls from 2014 to 2022, see [EBA, 2023, Annex I, p. 162] for more details and e.g. Thiele [2020] for further discussion on the asymmetry in financial data.
@@
$\widehat{\operatorname{\mathrm{VaR}}}_{1\%}(\mathbf{x}):=0.49x_{(2:250)}+0.51x_{(3:250)},$ (6.1)
@@
The numerical study highlights substantial differences between estimator performance across the non-overlapping and overlapping settings. In the non-overlapping framework (Table 4), estimators #–#3 generally provide the most accurate results, while estimators #4–#6 are clearly conservative, which is consistent with their Pareto tail based construction based on conservative simplifications. Among the former group, estimator #2 performs best with respect to fit metrics, whereas estimator #3 often delivers superior bias control and confidence coverage under heavy-tailed distributions. Overall, the ES estimators show a modest advantage over the benchmark $\operatorname{\mathrm{VaR}}$ estimator in terms of fit, but confidence threshold behavior is more sensitive to distributional assumptions; nevertheless, bias levels remain broadly in line with the traffic-light tolerance levels reported in Moldenhauer and Pitera [2019], where 0–5% CT output corresponds to the green performance zone and 5–10% CT output corresponds to the amber performance zone.
@@ Acknowledgements
Igor Cialenco acknowledges support from the US National Science Foundation grant DMS-2407549. Damian Jelito acknowledges support from the National Science Centre, Poland, via project 2024/53/B/ST1/00703. Marcin Pitera acknowledges support from the National Science Centre, Poland, via project 2024/53/B/HS4/00433. Martin Aichele thanks Carlo Acerbi for helpful discussions on preparatory work on ES estimation.

### v2

@@ Abstract
estimation of risk measures , coherent risk measure , coherent risk estimator , expected shortfall , value at risk
@@ 1 Introduction
In this work, we focus on the second step and adopt a novel perspective, fundamentally different from the existing literature, focused on the estimation function of $\rho$, say $\hat{\rho}$, which is later used to estimate $\hat{\rho}(X)$. We argue that, similar to the risk measures themselves, any estimation must also satisfy a set of desirable financial normative properties, postulated a priori, and in addition be a ‘good approximation’11endnote: 1There is a subtle distinction between estimation and approximation. Estimation refers to inferring an unknown true value from limited data, with an emphasis on its statistical properties. Approximation, in contrast, involves simplifying a known value or formula to make it computationally tractable, while analyzing the resulting numerical error. Our approach in this work combines elements of both, but for consistency we refer to them jointly as estimation.. This approach is motivated by the recognition that risk quantification procedures serve primarily to determine capital reserves for mitigating exposures, rather than to provide mere approximations of intrinsic risk values. To this end, we introduce the notion of the coherent risk estimator (CRE) that maps actual P&L samples to real numbers, that is monotone, cash additive, positive homogeneous, and subadditive when applied to a sample X.
@@
Our key theoretical contribution is the development of the robust representations of all CREs, see Theorem 4.1. Similar results are obtained for law-invariant22endnote: 2Similar to risk measures, a law-invariant CRE does not depend on the ordering of the input sample. See Definition 3.2. CREs, a sublcass of CREs. In fact, we prove that a CRE is law-invariant if and only if it can be represented as a supremum over a set of $L$-estimators; see Theorem 4.2. Moreover, assuming additionally that a CRE is also comonotonic, we show that it can be represented as an $L$-estimator; see Theorem 4.10. The importance of such results is evident. This approach provides a canonical framework for constructing risk estimators that are designed to possess the desired risk management properties. Furthermore, it enables not only the systematic selection of favorable estimators from the wide variety of existing alternatives, but also ensures that the chosen methods remain practically applicable, are integral, robust and distribution-free, and aligned with supervisory expectations, see e.g. ECB [2025]. Additionally, the strong connection between CREs and $L$-estimators facilitates the leverage of the extensive statistical literature about $L$-estimators and their various asymptotic properties, mostly applied to comonotonic and law-invariant CREs. Our approach also provides a fresh look on the robust representation theory for risk measures, in which the duality is build directly into the estimation formula rather than being imposed only on a theoretical risk metric level. Throughout the paper, we present various examples of CREs, with particular emphasis on estimators of $\operatorname{\mathrm{ES}}$ and their properties, including a dedicated section comparing different $\operatorname{\mathrm{ES}}$ estimators (see Section 6). We also discuss methods for constructing CREs for a given coherent risk measure using plug-in procedures, which recover estimators of several classical risk measures, including $\operatorname{\mathrm{ES}}$.
@@
Third, in the regulatory Internal Model Approach (IMA) for Pillar I bank models, the 10-day $\operatorname{\mathrm{VaR}}$ and 10-day $\operatorname{\mathrm{ES}}$ at the confidence levels $\alpha=1\%$ and $\alpha=2.5\%$, respectively, are the key reference market risk capital metrics, see BCBS [2006, 2019] for details.33endnote: 3In most practical applications the confidence level $\alpha$ is set to $1\%$, $2.5\%$, or $5\%$. Two main conventions are used when quoting confidence levels for $\operatorname{\mathrm{VaR}}$ and $\operatorname{\mathrm{ES}}$: the left-tail convention, adopted in this work and common in the risk measurement literature, and the right-tail convention, which reports thresholds of the form $1-\alpha$ (e.g., $99\%$, $97.5\%$, $95\%$) and is more widespread in risk management and regulatory practice, where the right tail represents losses. Although the risk measures themselves are fixed, regulatory and supervisory bodies rarely prescribe explicit estimation formulas. Two notable exceptions are the following locally implemented formulas used for Pillar I capital calculations: the minimal Stressed VaR formula under the Basel II framework [PRA, 2020, Article 10.2] and the EU Stress Scenario Risk Measure under the Basel III framework [EU, 2024b, Article 11]. More generally, regulatory and supervisory texts specify desired properties of estimators – such as conceptual soundness, proven backtesting track record, or distribution-free character – rather than their explicit form, see e.g. [ECB, 2025, Section 5.3], [PRA, 2020, Section 10], [CBUAE, 2023, Market Risk Standards & Risk Management Standards], or [HKMA, 2024, Section 4.5]. In practice, market risk estimators are often non-parametric and based on order statistics from historical simulation P&Ls, i.e., sorted P&L sample values. For $\operatorname{\mathrm{VaR}}$ and $\operatorname{\mathrm{ES}}$, the estimators are typically linear combinations of order statistics, with coefficients independent of the distribution of $X$; see EBA [2025] for typical VaR look-back period and weighting choices. These observations further motivates the relevance of the (comonotonic) CRE representations developed in this paper. We refer to Section 6 for a more detailed discussion of various nonparametric $\operatorname{\mathrm{ES}}$ estimations.
@@ 3 Coherent risk estimators
    Monotonicity, for any $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{n}$ such that44endnote: 4For vector order $\mathbf{x}\leq\mathbf{y}$ we use the component wise comparison $x_{i}\leq y_{i}$, for $i=1,2,\ldots,n$. $\mathbf{x}\geq\mathbf{x}^{\prime}$, we have $\hat{\rho}_{n}(\mathbf{x})\leq\hat{\rho}_{n}(\mathbf{x}^{\prime})$;
@@ 4.1 Comonotonic CREs and their representation as L-estimators
where $\xrightarrow{a.s.}$ stands for $\mathbb{P}$-almost sure convergence55endnote: 5In this work, for simplicity, we focus on $\mathbb{P}$-a.s. convergence of the estimators, which corresponds to strong consistency in classical statistics. Nevertheless, most of the results can be extended to weaker forms of convergence, such as convergence in probability (i.e., weak consistency)..
@@ 6 Numerical study: comparison of ES estimators based on L-statistics
Taking into account the nature of the considered metrics, we categorise the first two metrics (AE and SE) as Fit metrics, the next two (SB and RB) as Bias metrics, and the last one (CT) as Confidence metric. It should be noted that while our assessment is focused on estimators’ statistical properties, ES estimators are effectively point-forecast metrics focused on tail risk quantification. Thus, due to the risk-oriented nature of the ES estimators, the output of the AE and SE fit metrics should be handled with care. Those metrics may not be optimal for comparative accuracy assessment (and selection algorithms), since they are not elicitable for ES; see Gneiting [2011]. In particular, ES itself is not an elicitable risk measure, which makes comparative estimator analysis based on backtesting or scoring challenging. It is known that the joint metric $(\operatorname{\mathrm{VaR}},\operatorname{\mathrm{ES}})$ is elicitable, see Fissler and Ziegel [2016], but the associated scoring functions are difficult to compute and interpret. Moreover, the resulting comparisons are often statistically insignificant, may depend on the specific choice of scoring function, and require joint consideration of both ES and VaR estimation methodologies. Due to these challenges, we have decided to present, in addition to AR and SE, another set of metrics related to bias, as well as confidence measurement. We recall that the purpose of this paper is to provide a comparison of ES estimators based on $L$-statistics in the context of their CRE representation and the chosen scheme. A comprehensive assessment and comparative performance analysis are beyond the scope of this paper. For in-depth quantitative studies based on both simulated and market data, we refer to e.g. Chen [2008], Righi and Ceretta [2015], Bayer and Dimitriadis [2022], Barendse et al. [2023].
@@
In the next numerical example, we calculate the benchmark metrics for selected families of distributions using Monte Carlo simulations. For comparability, we consider the same family of distributions as in [EBA, 2023, Annex I]. This constitutes eight different distributions: standard normal distribution, Student’s $t$-distribution with $\nu=5$ degrees of freedom, and six different normal-inverse Gaussian (NIG) distributions with parameters $(\alpha,\beta)\in\{(0.4,0.14),(0.4,-0.14),(0.55,0.3025),(0.55,-0.3025),(0.4,0.22),(0.4,-0.22)\}$; the location parameter $\mu=0$ and the scale parameter $\delta=1$ are standardized. Skewness and kurtosis of the NIG distributions are known analytically and as NIG distributions are closed under convolution also for their rolling sums, [Scott et al., 2011, Section 6.1]. For the considered values of $(\alpha,\beta)$ parameters, the skewness and excess kurtosis values are $(0.9460,3.6282)$, (0.9460,3.6282)(-0.9460,3.6282), $(1.3601,4.5051)$, (1.3601,4.5051)(-1.3601,4.5051), $(1.8702,8.5174)$, and (1.8702,8.5174)(-1.8702,8.5174). The choice of $\alpha$ and $\beta$ aims to consider various signs and magnitudes of the skewness and the kurtosis comparable to trading book P&Ls from 2014 to 2022, see [EBA, 2023, Annex I, p. 162] for more details and e.g. Thiele [2020] for further discussion on the asymmetry in financial data.
@@
$\widehat{\operatorname{\mathrm{VaR}}}_{1\%}(\mathbf{x}):=-\left(0.49x_{(2:250)}+0.51x_{(3:250)}\right),$ (6.1)
@@
The numerical study highlights substantial differences between estimator performance across the non-overlapping and overlapping settings. In the non-overlapping framework (Table 4), estimators #1–#3 generally provide the most accurate results, while estimators #4–#6 are clearly conservative, which is consistent with their Pareto tail based construction based on conservative simplifications. Among the former group, estimator #2 performs best with respect to fit metrics, whereas estimator #3 often delivers superior bias control and confidence coverage under heavy-tailed distributions. Overall, the ES estimators show a modest advantage over the benchmark $\operatorname{\mathrm{VaR}}$ estimator in terms of fit, but confidence threshold behavior is more sensitive to distributional assumptions; nevertheless, bias levels remain broadly in line with the traffic-light tolerance levels reported in Moldenhauer and Pitera [2019], where 0–5% CT output corresponds to the green performance zone and 5–10% CT output corresponds to the amber performance zone.
@@
Funding
Igor Cialenco acknowledges support from the US National Science Foundation grant DMS-2407549. Damian Jelito acknowledges support from the National Science Centre, Poland, via project 2024/53/B/ST1/00703. Marcin Pitera acknowledges support from the National Science Centre, Poland, via project 2024/53/B/HS4/00433.
@@ Acknowledgements
Martin Aichele thanks Carlo Acerbi for helpful discussions on preparatory work on ES estimation.